D Question 7 Use the Pre-class table.xlsx document in Module Lecture 16, it contains molecular weight data for polytetrafluoroethylene. Calculate the number-average molecular weight. O 49,800 44,800 54,800 Question 8 D use the Pre-class table.xlsx document in Module Lecture 16, it contains molecular weight data for polytetrafluoroethylene. Calculate the weight-average molecular weight 55,200 o 50,200 60,200
media%2F37f%2F37fc3395-0f8f-4a0e-a5d1-3aCheck the answer and correct the wrong one
Thank you
Answer #1

Q7: The formula for number average molecular weight, Mn=?xiMi=x1M1+x2M2+....x8M8 =49800 g/mol. The calculation detail is given in the attached picture.

Q8: The formula for weight average molecular weight, Mw=?xiWi=x1W1+x2W2+....x8W8 =55200 g/mol. The calculation detail is given in the attached picture.

Q9:  Degree of polymerization=Dp=Mn/m. Here Mn is number average molecular weight and m is the molecular weight of polytetrafluoroethylene (C2F4)n.

m=(molecular weight of C)x2+ (molecular weight of F)x4 =12x2 + 19x4 =100 g/mol.

Therefore= Dp= 49800/100=498

1 Molecular Weight Range (g/mol) 2 10000-20000 3 20000-30000 Number average molecular weight ExiMi-49800 g/mol WI Mean Mi xiM